OpenStudy (anonymous):

solve the equation. 3x + 17 -5x = 12 - (6x + 3) need help with steps, not sure if im doing it right

OpenStudy (kamille):

alright, it is pretty straight forward. can you show what you have done so we can check if you did it alright?

OpenStudy (anonymous):

looking for x? First distribute the negative to the parentheses and get rid of them. then combine like terms.

OpenStudy (anonymous):

3x + 17 - 5x = 12 -6x -3 8x + 17 = 9 -6x 8 = -14x?

OpenStudy (anonymous):

are those steps right, cause i feel like theyre not

OpenStudy (kamille):

yea, all good, just some algebra mistakes...\[3x-5x \neq8x\] it should be like that: 3x+17-5x=12-(6x+3) -2x+17=12-6x-3 4x=-8 can you solve for x now?

OpenStudy (anonymous):

ohh okay so would x = -2

OpenStudy (kamille):

yes, that's correct! well done

OpenStudy (anonymous):

Okay thanks so much! :)